Hello out there,I had some problems that I need some help with badly,and thanks to whoever can help,the first question is,Determine if the equation is linear,and if it is put it in standard form.9y=4x-18.The next one is,xy=4.The next one is,x(with a 2 next it,I guess it means to the second power)-4=y.And the last problem is kinda hard,Two diced are rolled.Express a relation in a table for the possible numbers showing on the first die x and the second die y if the outcome is two odd numbers whose sum is a least 8.Thanks to whoever can help.
Answer by stanbon(75887)

the first question is,Determine if the equation is linear,and if it is put it in standard form.9y=4x-18.
It is linear because both x and y are to the 1st power
standard form is ax+by=c
For you 4x-9y=18 is the standard form
----------------
The next one is,xy=4.
This is not linear because the xy term is degree two.
--------------
The next one is,x(with a 2 next it,I guess it means to the second power)-4=y.
x^2-4=y
y=x^2-4
Not linear because the x is second power.
-----------
And the last problem is kinda hard,Two diced are rolled.Express a relation in a table for the possible numbers showing on the first die x and the second die y if the outcome is two odd numbers whose sum is at least 8.Thanks to whoever can help.
1 and anything will not add up to 8
2 and anything will not have two odd numbers
(3,5) works
4 and anything will not have two odd numbers
(5,3) works
6 and anything will not have two odd numbers
==========
Cheers,
Stan H.